SpaceRef

SpaceRef


NASA Hubble Space Telescope Daily Report # 3803

Status Report From: Space Telescope Science Institute
Posted: Wednesday, February 23, 2005

image

HUBBLE SPACE TELESCOPE - Continuing to collect World Class Science

DAILY REPORT        # 3803

PERIOD COVERED: DOY 53

OBSERVATIONS SCHEDULED

ACS/HRC/WFC 10444

ACS Two-gyro CVZ Test

4 orbits Only 1 visit. target = star cluster {CVZ star cluster??} Repeat first proposal visits 1-3. guide star V=13 goals: CVZ tracking, PSF, PSF repeatability, pointing stability {requirements 1a, 2, 4, 8} priority medium

ACS/WFC/HRC/NIC2 10443

ACS & WFPC2 Two-gyro PSF, pointing and dither test

12 orbits High priority observations {5 orbits} will be observed together at start of two- gyro SMS. Remaining 7 orbits provide further tests and additional experience in two-gyro mode. Goals: PSF, PSF repeatability, pointing and stability within the orbit {requirements 1a, 1c, 2 , 4}

FGS 10457

2-Gyro F2G Test Miscellaneous Orbits

This proposal is part of the February on-orbit 2-gyro testing.

FLIGHT OPERATIONS SUMMARY:

Significant Spacecraft Anomalies: (The following are preliminary reports of potential non-nominal performance that will be investigated.)

HSTAR 9717:  GSacq(2,3,2) failed to RGA Control.GSacq(2,3,2) scheduled for 053/09:53:53 - 10:02:48 Failed to RGA control due to Loss of Lock on FGS #3. This occurred after acquiring Fine Lock during F2G On-Orbit Test. No Observations affected. Under investigation.

HSTAR 9719:  FHST Obad Failure.FHST OBAD (But No Maneuver) scheduled for 054/05:11:36 Failed due to too many angle checks. A 486 STB message (1904) was received due to the failure. The Vehicle Axis Errors were 403.532, -168.097 and 239.083 (Arcsec). Under investigation.

HSTAR 9720:  GSacq(1,3,3) Failed to RGA Control.GSacq(1,3,3) scheduled for 054/06:40:46 - 06:48:54 Failed to RGA Control due to Search Radius Limit Exceeded on the Secondary FGS (#3). A Map was commanded at 06:48:01 and had an RSS value of 7.2 (arcsec). No observations affected.  Under investigation.

COMPLETED OPS REQs:

  • 17378-0 - Reset FGS1 K16ACLIM to 4,982 @ 053/1135z
  • 17370-0 - Continuous Engineering Coverage for F2G OOT (ESTR2 Record to 67.27%) @ 053/1223z-1554z
  • 17384-0 - Genslew for proposal 10445 - slot 1 @ 053/1557z
  • 17384-0 - Genslew for proposal 10445 - slot 2 @ 053/1559z
  • 17384-0 - Genslew for proposal 10445 - slot 3 @ 053/1600z
  • 17370-0 - Continuous Engineering Coverage for F2G OOT (ESTR2 to BOT) @ 053/2219z
  • 17370-0 - Continuous Engineering Coverage for F2G OOT (ESTR2 Record to 12.72%) @ 053/2334z-054/0011z
  • 17387-0 - Dump Tables 368, 369, 370 @ 054/0536z
  • 17388-0 - OBAD Command @ 054/0648z
  • 17379-1 - Reset FGS1 K16ACLIM to 4,982 @ 054/0917z

OPS NOTES EXECUTED: 1307-0 - TGS OOT Signal Strength Monitor (in progress)

                           SCHEDULED     SUCCESSFUL    FAILURE TIMES 
 FGS 
 Gsacq                 5                        3                053/0953z 
 (HSTAR 9717); 
                                                                                054/0640z 
 (HSTAR 9720) 
 FGS Reacq                 5                        5 
 FHST Update 
 OBAD                        19                      18 
 054/0511z (HSTAR 9719) 
 LOSS of LOCK 

SIGNIFICANT EVENTS:

Continuing execution of Two-Gyro Science on-orbit test.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks